OpenStudy (anonymous):

Solve the system. x – 2y = 10 3x + 2y = 14 (–8, 6) (6, –8) (6, –2)

OpenStudy (igreen):

We can add the two equations. x - 2y = 10 3x + 2y = 14

OpenStudy (igreen):

x + 3x = ? -2y + 2y = ? 10 + 14 = ?

OpenStudy (igreen):

Can you add those?

OpenStudy (igreen):

x + 3x is the same as 1 + 3, but we put an 'x' next to it.

OpenStudy (igreen):

So what's 1 + 3?

OpenStudy (anonymous):

I got it thanks